The vapor pressure of solid iodine is: p.7.632 Torr @ T-68.8 ℃ P-615.7 Torr @ T-154.3 ℃ Determine Δ for 12(s).
Answer #1

P1   = 7.732 torr              T1 = 68.8+273    = 341.8K

P2   = 615.7 torr                T2   = 154.3+273 = 427.3K

logP2/P1    = \DeltaH/2.303R [1/T1 -1/T2]

log615.7/7.732   = \DeltaH/2.303*8.314 [1/341.8-1/427.3]

1.9010                   =\DeltaH(0.00292-0.00234)/19.147

1.9010                  = \DeltaH*0.00058/19.147

\DeltaH                      = 1.9010*19.147/0.00058    = 62755.943J      = 62.756KJ >>>>>answer